Fermi National Laboratory
 
Operation Summary for February 13, 2012 - February 15, 2012
 

Monday February 13

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Water Group personnel reported at 8:21 AM that they were making up water for the MI magnet LCW system. FESS personnel reported at 8:22 AM that they were taking down feeder 46A for a week of maintenance. Operators loaded the Recycler with protons at 8:54 AM. A Linac person adjusted the low-energy collimator at 9:54 AM. A MI person began a 30-minute study at 10:03 AM. At 10:26 PM, a Linac quadrupole magnet power supply (L:QPS108) tripped off; it reset. Operators prepared Meson enclosures M03 and M04 for access at 12:46 PM. A MI person began a g-2 (g minus 2) study at 1:15 PM. Operators prepared Meson enclosures M01 and M05 for access at 1:23 PM. A Booster person began a 10-minute study at 2:08 PM. At 2:22 AM, the Booster West Anode Power Supply tripped off due to a Booster RF station (BRF16); Booster resumed sending beam at 2:42 PM. At 2:54 PM, a repeater went down; Controls personnel investigated and found an Uninterruptable Power Supply (UPS) had failed. They replaced the UPS. The Meson enclosures M01, 3, 4, and 5 were all out of access by 3:35 PM.

During the evening shift, a Linac person restored the low-energy collimator to its original position at 4:46 PM. An MI person completed his g-2 study at 6:10 PM. At 11:08 PM, a Linac RF station (LRF4) tripped off; it reset.

Tuesday February 14

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:52 AM, LRF4 began having gradient problems; Operators investigated and reported the problem was due to the driver crowbar bit and at 2:10 AM, Operators called in a Linac technician. At 2:18 AM, Operators reported that LRF4 started working normally. At 5:59 AM, LRF4 tripped off; it reset. Operators reported at 6:20 AM, that there had been a number of intermittent MiniBooNE beam permit trips due to bad beam positions; Operators adjusted the Booster frequency curve delay, which resolved the problem. At 7:59 AM, Neutrino Therapy Facility personnel prepared to treat two patients later in the morning.

During the day shift, at 8:09 AM, LRF4 tripped off; it reset. Water Group personnel reported at 8:14 AM that they were regenerating column #2 at CUB. Operators prepared Meson enclosures M03 and M04 for access at 8:30 AM. At 8:33 AM, a Booster vacuum pump (B:IP52) tripped off; it reset. At 10:15 AM, a Recycler kicker (R:KPR4A) tripped on a pre-fire; it reset. At 10:15 AM, BRF11 tripped off on a ground fault indication; it reset. Operators searched and secured the Neutrino Area beamline G2 enclosure at 11:12 AM. Operators reported at 11:18 AM that a MI kicker (I:KPS6NF) tripped off and caused the NuMI beam permit to also trip; everything reset. LRF4 tripped off at 12:41 PM; Linac technicians investigated and replaced its interlock module. Linac resumed sending beam at 1:08 PM. A Booster person began a study at 1:12 PM. Operators searched and secured the MTest enclosures at 1:49 PM. Operations established beam to MTA at 2:47 PM. Operation established beam to MTest for the “EDIT School” at 2:59 PM, which will run during weekdays for the next two weeks. EDIT stands for Excellence in Detector and Instrumentation Technology. BRF19 tripped off at 3:19 PM; it reset.

During the evening shift, Operations disabled beam to MTA at 4:57 PM. The Booster studies ended at 4:58 PM. LRF4 tripped off for the first of three times in nine minutes; Linac personnel investigated. At 5:39 PM, L:QPS304 tripped off and would not reset; Linac personnel investigated. The MTest EDIT School ended for the day at 6:04 PM. Linac resolved its RF and quadrupole problems and resumed sending beam at 7:34 PM. A returning Meson FTBF experiment T-979 began taking beam. Operations reported that the output of the H- Source had dropped off since beam returned; Linac personnel had the I- Source up a running by 8:27 PM. At 10:31 PM, BRF11 had tripped off for the first of four times in about an hour; it reset. Operations disabled the T-979 beam at 10:37 PM.

Wednesday February 15

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:15 AM, BRF11 tripped off for the first of four times in 53 minutes; it reset. Operators who were investigating the BRF11 ground fault trips found that sparking was the real cause of the trips; they lowered the station’s output, which resolved the sparking problem. Operators completed a Booster injection tune at 3:39 AM. Operations changed the MI chromaticity at 4:30 AM. At 6:13 AM, the I- Source sparked and caused L:QPS104 to trip off; it reset.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo wants the complex to operate with as few beam disruptions as possible for the EDIT School.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/15/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory